# CHARTOM: A Visual Theory-of-Mind Benchmark for Multimodal Large Language Models

### v1

@@ Abstract
We introduce CHARTOM, a visual theory-of-mind benchmark for multimodal large language models. CHARTOM consists of specially designed data visualizing charts. Given a chart, a language model needs to not only correctly comprehend the chart (the FACT question) but also judge if the chart will be misleading to a human reader (the MIND question). Both questions have significant societal benefits. We detail the construction of the CHARTOM benchmark including its calibration on human performance.
@@ 1 Introduction
1.1 Background on Theory of Mind
For AI to better assist humans, it must know not just factual truth but also how humans perceive the truth. The two could differ for many reasons such as information asymmetry or cognitive ability. If AI can detect the difference, it can help humans accordingly. The ability for an individual to reason about how others think (rather than the factual truth) is known as theory of mind [30]. A classic example is the Sally-Anne test [38, 2]: Sally hides a marble in a basket then leaves the room. Anne moves the marble to a box while Sally is away. Sally comes back. At this point, one can ask an observer two types of questions:
@@ 1.1 Background on Theory of Mind
1.2 Background on Data Visualizing Charts
Our theory of mind benchmark tests visual perception instead. Specifically, our focus is on misleading data visualizing charts, e.g. bar charts, line graphs, pie charts etc. Charts are widely used for conveying quantitative information to human readers. They lend credibility to accompanying messages and hence have a strong impact on human decision making. On the flip side, however, if charts are misleading, they can amplify the impact of misinformation [22, 29]. Figure 1 reproduces a misleading chart whose $y$-axis is inverted, which may give the impression that enacting Florida’s “stand your ground” law reduced gun deaths. Misleading charts are unfortunately prevalent in many areas of our lives. For example, a review of medical advertisements found that 1/3 of charts were misleading [7]. A review of charts in Korean news about the COVID-19 pandemic showed that about 30% of bar charts and 44% of pictorial charts were misleading [21]. A U.S.-based analysis of visuals that had been identified as misleading by human fact checkers showed that more than half of these visuals were used as evidence for fake news about COVID-19 [4]. Because journalists often use charts to make narratives appear objective and to engage viewers emotionally [32], misleading charts can make an inaccurate story more deceptive [8, 37].
@@ 1.2 Background on Data Visualizing Charts
Existing computer vision research in charts largely aimed at chart comprehension. Classic approaches often employ an intermediate step of chart-to-data parsing, then use the data for downstream tasks such as visual question-answering, reasoning and summarization. To this end, [5, 28] applied deep learning based object detection methods to extract data, while others [1, 6, 26, 16, 34] applied semi-automated off-the-self OCR text recognition systems [3, 17]. Newer MLLM-based approaches tends to be end-to-end, such as training the model on extensive pre-training datasets [25, 12, 27, 40], few-shot tuning methods which fine-tune pre-trained models with a small number of task-specific examples [9, 23, 19], and even zero-shot learning where the model makes predictions on tasks it has not been explicitly trained on [39, 15]. The newer approaches are performant, though occasionally still suffer from hallucination in data extraction and reasoning [14, 13].
However, we point out a significant shortcoming in chart comprehension: many misleading charts in fact contain the correct data. Take Figure 1 for example. Despite the unusual flipped $y$-axis, a meticulous reader (or AI) can indeed recover all the correct (year, deaths) data points from the chart. Therefore, ironically, a perfect chart comprehension system is not directly suited for judging how misleading a chart may be to humans. This is the same distinction between factual and theory of mind questions. We can ask a FACT question: What data is objectively presented in the chart? We can also ask a MIND question: What will a typical human reader perceive from the chart? Chart comprehension systems aim to solve the FACT question. To the best of our knowledge, the present paper is the first to address the MIND question in charts instead.
1.3 Our Contributions
Our contribution is to create a visual theory-of-mind benchmark CHARTOM (for CHARt Theory of Mind) consisting of specially designed charts. The benchmark covers common varieties of chart types, and we planted representative visual manipulations into the charts. The key to the benchmark is to obtain MIND ground truth, i.e. “degree of misleadingness to humans”, for each chart. This ground truth enables future benchmark users to judge whether their AI correctly answers the MIND questions. We obtained the MIND ground truth via human experiments.
@@ 2 The CHARTOM Benchmark
(a) original bar chart G7_Q1_1
(b) manipulated bar chart G7_Q1_2
Figure 2: A pair of example charts in CHARTOM
In this section we present our visual theory of mind benchmark CHARTOM. We first detail our design of charts to accommodate both FACT and MIND type of questions in the benchmark, then explain how we obtain the MIND ground truth with human experiments.
@@ 2.1 Design of the CHARTOM Benchmark
An example. The pair of charts in Figure 2 share the same underlying data: (year=2018, dogs=250) and (year=2019, dogs=300). The left chart is standard, while we intentionally planted a potentially misleading manipulation to the right chart: we truncated the $y$-axis range so it starts at 240 instead of 0. This is to create a perceptual effect: the 2019 bar now appears much taller than the 2018 bar. For each of the two charts, we ask the following FACT and MIND questions.
@@
    What fraction of typical university students do you predict will be misled by the chart when answering the question? First give your prediction as a decimal number between 0 and 1, then justify your prediction in words.
Answers to the FACT question is easy to judge. It is important to note that in both charts by design we ensured the underlying data can be exactly recovered, regardless of the manipulation. Therefore, a perfect MLLM-based AI, with proper reasoning, can always correctly answer the FACT question on both charts: the answer is $\frac{1}{250}\times 300=1.2$ million bags.
In contrast, the MIND question is trickier to judge. On the left chart where the design is standard, one might still argue that the chart is not perfect and could be a little bit misleading to humans: the numbers are too small, there is no grid guide lines to help visual alignment, etc. Conversely, on the right chart where we intentionally planted the truncated $y$-axis because we presumed it to be misleading based on existing psychology literature (e.g. [24]), we do not know a priori the actual effect on human readers. One way to obtain ground truth on the MIND question is to actually conduct human experiments, which we explain in section 2.2.
The CHARTOM benchmark. Our entire CHARTOM benchmark follows the same design principle. The benchmark consists of 112 charts, covering five prevalent chart types: line, bar, pie, scatter plot and map. The charts are manually designed to be visually tidy. We present selected charts from the benchmark in Table 1 and Appendix A.
The 112 charts come in 56 pairs (one pair is already shown in Figure 2). Each pair has an original (suffix _1) and a manipulated (suffix _2) version, the latter incorporates one of many visualization fallacies suggested by the psychology literature, see the ‘manipulation’ column in Table 1. The manipulated version is intended to confuse casual human readers. Each pair of charts also comes with a FACT question and a MIND question. The FACT question is typically a simple word problem based on chart comprehension, see examples in Appendix A. There are three types of FACT questions: multiple choices, free text entry, and sorting. Importantly, the true underlying data can still be recovered from either version if one is careful. This means that the FACT question has the same answer on both versions. The MIND question has the same format as shown earlier for all charts.
@@
  • 112 charts: 56 pairs of original and manipulated versions.
@@
  • Each chart comes with a MIND question, and the answer key is the HMI estimated from human experiments (see Section 2.2).
To avoid web crawlers inadvertently adding the benchmark to future MLLM training data, the CHARTOM benchmark is not online but is available upon request. Please contact Prof. Jerry Zhu at jerryzhu@cs.wisc.edu.
original manipulated
chart type ID HMI ID manipulation HMI
line G4_Q1_1 0.03 G4_Q1_2 inconsistent x-axis 0.97
line G3_Q1_1 0.03 G3_Q1_2 inverted y-axis 0.81
line G9_Q1_1 0.06 G9_Q1_2 dual axes 0.74
line G1_Q1_1 0.14 G1_Q1_2 truncated y-axis 0.40
line G2_Q1_1 0.07 G2_Q1_2 compressed y-axis 0.04
bar G12_Q1_1 0.28 G12_Q1_2 3D effect 0.96
bar G7_Q1_1 0.18 G7_Q1_2 truncated y-axis 0.29
bar G8_Q1_1 0.09 G8_Q1_2 compressed y-axis 0.06
bar G10_Q1_1 0.02 G10_Q1_2 pictorial bars 0.06
scatter G6_Q1_1 0.03 G6_Q1_2 swapped axis 0.71
scatter G5_Q1_1 0.02 G5_Q1_2 logarithmic y-axis 0.63
map G15_Q1_1 0.01 G15_Q1_2 inverted color scale 0.10
pie G13_Q1_1 0.26 G13_Q1_2 3D effect 0.90
pie G14_Q1_1 0.66 G14_Q1_2 3D effect + pop out 0.99
Table 1: Selected chart pairs in the CHARTOM benchmark. Each row is for a pair of original and manipulated charts. HMI standards for Human Misleadingness Index, a number between 0 and 1 where a larger value suggests the chart is more misleading to humans.
2.2 HMI: Ground Truth for MIND Questions
A key part of our benchmark is the ground truth answer for the MIND question on each chart, namely whether that chart is misleading to typical human readers.
2.2.1 Estimating HMI from Human Experiments
We propose Human Misleadingness Index (HMI), a number between 0 and 1 for each chart where a larger value means the chart is more misleading to humans. HMI is estimated with the following process.
We conducted a human experiment on 68 university students with IRB approval. We did not directly ask human participants the MIND question because we do not anticipate people to be able to give reliable answers—after all, many were themselves misled by some of our charts. Instead, we asked them the FACT question, which is designed so that users who are misled by the chart are likely to give wrong answer. Subsequently, we use the performance of users on the FACT question to establish the ground truth for the MIND question. Concretely, the HMI for a chart is defined as the percent of our study population who did not give an acceptable answer on the FACT question. We will define acceptability in the next section. We argue that HMI is a surrogate measure for how misleading the chart is. This definition is motivated by two factors:
@@ 2.2.1 Estimating HMI from Human Experiments
    The arithmetic involved in solving the FACT questions is straightforward for the education level of our participants. Indeed on most original (not manipulated) charts the HMI is small.
@@
    On many manipulated charts the HMI is higher compared to their original chart counterpart. Since the only difference between the pair is the planted manipulation in data visualization, we can attribute the increase in HMI to the manipulation being more misleading to humans.  [^1] We randomized the order of original and manipulated charts in each pair when presenting them to human participants to minimize the order effect.
Our human experiments are described in greater detail in [31].
2.2.2 Defining Acceptable Human Answers
Recall the FACT questions can be either multiple choices, sorting, or free text entry. It is straightforward to judge human answers for multiple choice and sorting questions (where we require all items are in the correct order), but it is more subtle for free text entry. For the purpose of calculating HMI, on free text entries we allow approximately correct answers. For example, we already discussed that the FACT answer is 1.2 for Figure 2. However, many participants understandably answered other numbers around 1.2, and a few participants answered drastically different numbers, see Table 2. A few remarks are in order:
  • For both the original (G7-Q1-1) and manipulated (G7-Q1-2) charts, the mode of the answer distribution is at the correct FACT answer 1.2.
  • For the manipulated chart G7-Q1-2, many participant answered 6. This answer was not observed for the original chart. We believe those participants were misled by the vertical perceptual bar heights. In other words, they used the wrong formula 1(1 vertical unit)×(6 vertical $\frac{1}{(1\mbox{ vertical unit})}\times(6\mbox{ vertical units})=6$ million bags.
  • The answer 300 is also frequent. We believe those participants simply read off the $y$-axis number of the 2019 bar.
  • Many participants answered other numbers around 1.2, 6, or 300.
original chart G7-Q1-1 answer 1 1.01 1.1 1.15 1.16 1.2 1.25 1.3 1.5 2 3 250 300 345 count 5 1 5 1 1 24 5 5 6 3 2 1 8 1 manipulated chart G7-Q1-2 answer 1 1.1 1.13 1.1666 1.2 1.25 1.26 1.3 1.5 5 6 30 56 300 count 4 2 1 1 22 4 1 2 11 4 7 1 1 7
Table 2: All participant answers on the FACT question for the pair of charts in Figure 2
We encountered similar diversity of human answers on other free text entry questions. These answers are indicative of the diverse cognitive processes in our participants, and are of independent interest to psychologists. From a computational perspective, we simplified the cognitive processes to define acceptable answers and assume that a participant will use one of following three problem solving strategies :
@@ 2.2.2 Defining Acceptable Human Answers
    The correct strategy. This leads to the correct FACT answer. For G7-Q1-1 and G7-Q1-2 the answer would be 1.2.
@@
    The misled strategy where they “fall for the manipulation”. For G7-Q1-2 the answer would be 6.

### v2

@@ Abstract
We introduce CHARTOM, a visual theory-of-mind benchmark for multimodal large language models. CHARTOM consists of specially designed data visualizing charts. Given a chart, a language model needs to not only correctly comprehend the chart (the FACT question) but also judge if the chart will be misleading to a human reader (the MIND question). Both questions have significant societal benefits. We detail the construction of the CHARTOM benchmark including its calibration on human performance. We benchmarked leading LLMs as of late 2024—including GPT, Claude, Gemini, Qwen, Llama, and Llava—on the CHARTOM dataset and found that our benchmark was challenging to all of them, suggesting room for future large language models to improve.
@@ 1 Introduction
For AI to better assist humans, it must know not just factual truth but also how humans perceive the truth. The two could differ for many reasons such as information asymmetry or cognitive ability. If AI can detect the difference, it can help humans accordingly. The ability of an individual to reason about how others think (rather than the factual truth) is known as the theory of mind [34]. A classic example is the Sally-Anne test [43, 3]: Sally hides a marble in a basket and then leaves the room. Anne moves the marble to a box while Sally is away. Sally comes back. At this point, one can ask an observer two types of questions:
@@
Our theory of mind benchmark tests visual perception instead. Specifically, our focus is on misleading data visualizing charts, e.g. bar charts, line graphs, pie charts, etc. Charts are widely used for conveying quantitative information to human readers. They lend credibility to accompanying messages and hence have a strong impact on human decision-making. On the flip side, however, if charts are misleading, they can amplify the impact of misinformation [24, 33]. Figure 1 reproduces a misleading chart whose $y$-axis is inverted, which may give the impression that enacting Florida’s “stand your ground” law reduced gun deaths. Misleading charts are unfortunately prevalent in many areas of our lives. For example, a review of medical advertisements found that 1/3 of charts were misleading [8]. A review of charts in Korean news about the COVID-19 pandemic showed that about 30% of bar charts and 44% of pictorial charts were misleading [23]. A U.S.-based analysis of visuals that had been identified as misleading by human fact-checkers showed that more than half of these visuals were used as evidence for fake news about COVID-19 [5]. Because journalists often use charts to make narratives appear objective and to engage viewers emotionally [36], misleading charts can make an inaccurate story more deceptive [9, 42].
@@
1.1 Related Works
Existing computer vision research in charts have largely focused on chart comprehension. Classic approaches often employ an intermediate step of chart-to-data parsing, then use the data for downstream tasks such as visual question-answering, reasoning, and summarization. To this end, [6, 31] applied deep learning-based object detection methods to extract data, while others [1, 7, 29, 18, 38] applied semi-automated off-the-self OCR text recognition systems [4, 19]. Newer LLM-based approaches tend to be end-to-end, such as training the model on extensive pre-training datasets [28, 14, 30, 46], few-shot tuning methods which fine-tune pre-trained models with a small number of task-specific examples [10, 26, 21], and even zero-shot learning where the model makes predictions on tasks it has not been explicitly trained on [44, 17]. The newer approaches are performant, though occasionally still suffer from hallucination in data extraction and reasoning [16, 15].
However, we point out a significant shortcoming in chart comprehension: many misleading charts in fact contain the correct data. Take Figure 1 for example. Despite the unusual flipped $y$-axis, a meticulous reader (or AI) can indeed recover all the correct (year, deaths) data points from the chart. Therefore, ironically, a perfect chart comprehension system is not directly suited for judging how misleading a chart may be to humans. This is the same distinction between factual and theory of mind questions. We can ask:
  • A FACT question: What data is objectively presented in the chart?
  • A MIND question: What will a typical human reader perceive from the chart?
Current chart comprehension systems aim to solve the FACT question. To the best of our knowledge, the present paper is the first to address the MIND question in charts instead.
Our Contributions:
Towards this end, we make the following contributions:
  1. 1.
    We create a visual theory-of-mind benchmark CHARTOM (for CHARt Theory Of Mind) consisting of specially designed charts. The benchmark covers common variety of chart types and planted visual manipulations in those charts as shown in Table 1
  2. 2.
    We designed two types of questions for each chart: 1.) FACT question that are based on factual information present in charts, and, 2.) MIND question based on human theory of mind.
  3. 3.
    The key to the benchmark is to obtain MIND ground truth, i.e. “degree of misleadingness to humans”, for each chart that has been obtained via human experiments. This ground truth enables accessing whether an AI correctly answers the MIND questions.
  4. 4.
    We test the performance of leading LLMs in Claude, Gemini, GPT, Llama, LLava, and Qwen series as of Nov 2024 and report our findings. We observe that all LLMs perform poorly on both FACT and MIND questions indicating a big scope of improvement.
@@ 2 The CHARTOM Benchmark
Figure 2: A pair of non-manipulated(left) and manipulated(right) charts in CHARTOM with truncated y-axis manipulation.
In this section, we present our visual theory of mind benchmark CHARTOM. We first detail our design of charts to accommodate both FACT and MIND type of questions in the benchmark, then explain how we obtain the MIND ground truth with human experiments.
@@ 2.1 Design of the CHARTOM Benchmark
An example. The pair of charts in Figure 2 share the same underlying data: (year=2018, dogs=250) and (year=2019, dogs=300). The left chart has a standard design which we call “non-manipulated”, while in the right chart we intentionally planted a potentially misleading manipulation: we truncated the $y$-axis range so it starts at 240 instead of 0. This is to create a perceptual effect: the 2019 bar now appears much taller than the 2018 bar. For each of the two charts, we ask the following FACT and MIND questions.
@@
    What fraction of typical university students do you predict will be misled by the chart when answering the question? First, give your prediction as a decimal number between 0 and 1, then justify your prediction in words.
The CHARTOM benchmark. Our entire CHARTOM benchmark follows the same design principle. The benchmark consists of 112 charts, covering five prevalent chart types: line, bar, pie, scatter plot, and map. The charts are manually designed to be visually tidy. We present selected charts from the benchmark in Table 1 and Appendix A.
The 112 charts come in 56 pairs (one pair is already shown in Figure 2). Each pair has a non-manipulated (suffix _1) and a manipulated (suffix _2) version, the latter incorporates one of many visualization fallacies suggested by the psychology literature, see the ‘manipulation’ column in Table 1. The manipulated version is intended to mislead casual human readers. Each pair of charts also comes with a FACT question and a MIND question. The FACT question is typically a simple word problem based on chart comprehension, see examples in Appendix A. There are three types of FACT questions: multiple-choice, free text entry, and sorting. Importantly, the true underlying data can still be recovered from either version if one is careful. This means that the FACT question has the same answer on both versions. The MIND question has the same format as shown earlier for all charts.
@@
  • 112 charts: 56 pairs of non-manipulated and manipulated versions.
@@
  • Each chart comes with a MIND question, and the answer key is the Human Misleadingness Index (HMI, see Section 3.2) observed in human experiments.
To avoid web crawlers inadvertently adding the benchmark to future LLM training data, the CHARTOM benchmark is not online but is available upon request.
Non-Manipulated Manipulated
chart type group ID HMI manipulation type HMI
line G1 0.51 inconsistent x-axis 0.68
line G2 0.76 compressed y-axis 0.88
line G3 0.03 inverted y-axis 0.83
line G4 0.05 truncated y-axis 0.92
line G9 0.19 dual axes 0.84
bar G7 0.31 truncated y-axis 0.53
bar G8 0.76 compressed y-axis 0.69
bar G10 0.19 pictorial bars 0.27
bar G12 0.11 3D effect 0.48
scatter G5 0.02 logarithmic y-axis 0.63
scatter G6 0.02 swapped axis 0.73
map G15 0.02 inverted color scale 0.15
pie G13 0.25 3D effect 0.79
pie G14 0.30 3D effect + pop out 0.79
Table 1: MIND ground truth answers (HMI) score averaged across various manipulation groups in the CHARTOM benchmark. Each row is for non-manipulated and manipulated version of charts within a manipulation group. HMI stands for Human Misleadingness Index, a number between 0 and 1 where a larger value suggests the chart is more misleading to humans.
3 Obtaining Ground Truth Answers
3.1 Ground Truth for FACT Questions
The FACT questions are based on ground truth data. Their ground truth answers are easily computable. Recall both non-manipulated and manipulated versions by design have the same underlying data, and that data can be exactly recovered regardless of visual manipulation. Therefore, a perfect human or AI with proper reasoning can always obtain the same correct FACT answer on both chart versions. For example, in Figure 2 the correct FACT answer is $1.21\times\frac{300}{250}=$ 1.2 million bags.
Specifically, on FACT questions we suggest using accuracy, defined as the number of correct answers over the number of FACT questions. To define correctness, we consider the three types of FACT questions separately:
@@ 3.1 Ground Truth for FACT Questions
    Multiple choice questions. An answer is correct if it precisely matches the answer key.
@@
    Ranking questions. A ranking answer is correct only if all items are ranked correctly.
  3. 3.
    Numerical (free text) questions. For numerical answers, the answer key provides the exact answer assuming perfect perception. But an exact match can be too strong an assumption for either humans or LLMs when they can enter any free-text number. To judge numerical answers, we suggest using a ±$\pm 10\%$ tolerance interval: if the answer key is $a$, then an LLM answer in the interval $[0.9a,1.1a]$ is deemed correct.
3.2 Ground Truth for MIND Questions
In contrast, the MIND question is trickier to judge. On the non-misleading chart in Figure 2(left) where the design is standard, one might still argue that the chart is not perfect and could be a little bit misleading to humans: the numbers are too small, there is no grid guidelines to help visual alignment, etc. Conversely, on the misleading chart (right) even if we intentionally planted the truncated $y$-axis because we presumed it to be misleading based on existing psychology literature (e.g. [27]), we do not know a priori the actual effect on human readers. To avoid these difficulties we propose the Human Misleadingness Index (HMI), a number between 0 and 1 for each chart which is the fraction of the human population that is misled by the chart. The correct answer to a MIND question is the HMI value.
To obtain HMI we conducted human experiments, which is expensive and time consuming but is the only proper method. Our human experiment involves 68 university students with IRB approval. We did not directly ask human participants the MIND question because we do not anticipate people to be able to give reliable answers—after all, many were themselves misled by some of our charts. Instead, we asked them the FACT question, which is designed so that users who are misled by the chart are likely to give wrong answers. We randomized the order of non-manipulated and manipulated charts in each pair when presenting them to human participants to minimize the order effect. We estimate the HMI of a chart by the percent of our study population who did not give a correct answer on the FACT question.
We argue that HMI is a surrogate measure for how misleading the chart is. This definition is motivated by two factors:
@@ 3.2 Ground Truth for MIND Questions
    The arithmetic involved in solving the FACT questions is straightforward for the education level of our participants. Indeed on most non-manipulated charts, the HMI is small.
@@
    On many manipulated charts the HMI is higher compared to their non-manipulated chart counterpart. Since the only difference between the pair is the planted manipulation in data visualization, we can attribute the increase in HMI to the manipulation being more misleading to humans.
Figure 3 shows the HMI values in the CHARTOM benchmark. These values are sorted from small (not misleading to humans) to large (very misleading). In addition, each HMI value is color-coded to indicate whether the chart is the non-manipulated version (blue) or the manipulated version (red). Generally speaking, the non-manipulated charts have small HMI, and the manipulated ones have large HMI. But this depends on the type of manipulation. As the HMI values in Table 1 indicate, some manipulations greatly misled people: inconsistent x-axis, inverted y-axis, dual axes plots, 3D effect, etc. While interestingly, some other manipulations did not confuse people much: compressed y-axis and pictorial bars are benign. Such differences are of interest to cognitive science. Our human experiments are described in greater detail in [35].
4 Benchmarking Leading LLMs
We evaluated leading LLMs including Claude Sonnet 3.5 [2], Gemini 1.5 Pro [40], GPT-4o [32], LLaVA-OneVision [25], Llama-3.2-90B [13], Qwen2-VL-72B [45] on both FACT and MIND questions and present our findings below. When evaluating LLMs, we suggest the following criteria:
  • On FACT questions, we suggest reporting accuracy.
  • On the numerical part of the MIND questions, we suggest reporting mean absolute error: $\frac{1}{n}\sum_{i=1}^{n}|m_{i}-a_{i}|$ where $n$ is the number of MIND questions, $m_{i}$ is an LLM’s answer on the $i$th MIND question, and $a_{i}$ is the HMI answer key to that MIND question.
  • On the justification part of the MIND questions, we suggest qualitative comparisons to the “chart manipulations” file provided in the benchmark. That file records the single manipulation we planted to the manipulated (suffix _2) version of each chart pair. If a manipulated chart has a high HMI value, a capable LLM should identify the corresponding manipulation as an important cause of misleadingness.
The following are our main findings.
4.1 Like Humans, LLMs are Fooled by Manipulations
We asked LLMs to answer FACT questions on both non-manipulated and manipulated charts and record their response. Figure 4 shows the error rate of each LLM separately on non-manipulated charts (blue), manipulated charts (orange), and all charts (green). We make the following observations:
  1. 1.
    LLMs perform significantly worse on manipulated charts as compared to non-manipulated ones. This is evident from the difference in height of blue and orange bars for each LLM in Figure 4. We performed a $\mathcal{X}^{2}$ test that achieved a p-value $\leq 0.05$ for all LLMs. This suggests that, like humans, LLMs are fooled by manipulations.
  2. 2.
    Top LLMs show near or superhuman performance on FACT questions. We observe that top LLMs like Claude, Qwen, and GPT-4o have performed better or very close to humans (on average) while others like Llama, Gemini, and LLaVA have shown relatively poor performance.